OpenStudy (idealist10):

Find solutions y1 and y2 of the equation y"=0 that satisfy the initial conditions y1(xo)=1, y'1(xo)=0 and y2(xo)=0, y'2(xo)=1.

OpenStudy (idealist10):

r^2=0 r=0 (double root) y=c1+c2x Now what? @SithsAndGiggles

OpenStudy (anonymous):

What are these initial conditions? Just general \(y(x_0)=0\) ?

OpenStudy (anonymous):

1, not 0*

OpenStudy (idealist10):

The initial conditions are\[y _{1}(x _{0})=1, y'_{1}(x _{0})=0, y _{2}(x _{0})=0, y'_{2}(x _{0})=1\]

OpenStudy (anonymous):

Alright I was just wondering if you were given a particular value for \(x_0\). From your solution we know the two fundamental solutions. \(y=C_1+C_2x\) admits \(y_1=1\) and \(y_2=x\) (or possibly the other way around, with \(y_1=x\) and \(y_2=1\) - we'll find out shortly). With \(y_1(x_0)=1\) and \({y_1}'(x_0)=0\), it's likely that \(y_1\) is the constant solution, since \[\begin{cases}y_1=C_1\\{y_1}'=0\end{cases}\quad\text{and}\quad\begin{cases}y_1(x_0)=1\\{y_1}'(x_0)=0\end{cases}~~\implies~~C_1=1\] Similarly, if \(y_2=C_2x\), then \[\begin{cases}y_2=C_2x\\{y_2}'=C_2\end{cases}\quad\text{and}\quad\begin{cases}y_2(x_0)=0\\{y_2}'(x_0)=1\end{cases}~~\implies~~C_2=1\] (and also that \(x_0=0\)).

OpenStudy (idealist10):

The answers are \[y _{1}=1, y _{2}=x-x _{0}\]

OpenStudy (idealist10):

How did you find C1=1 and C2=1 and also that xo=0?

OpenStudy (anonymous):

Ignore everything past "Similarly,..." That's wrong. I thought the work suggested that \(x_0=0\), but that's not the case, and it is in fact a general constant. If you can accept that the fundamental solutions are \(y_1=1\) and \(y_2=x\), finding the values of \(C_1\) and \(C_2\) is a matter of solving these simple IVPs. We have the ODE \(y''=0\) with initial conditions \(y_1(x_0)=1\) and \({y_1}'(x_0)=0\). You know the general solution is \(y=C_1+C_2x=C_1y_1+C_2y_2\). The principle of superposition allows us to consider one solution at a time. The first initial condition tells us that \[y_1(x_0)=1~~\implies~~C_1=1\] The second says \[{y_1}'(x_0)=0,\quad\text{which doesn't give us any more info}\]

OpenStudy (anonymous):

Now consider the second solution, \(y_2=C_2x\), and its initial conditions: \(y_2(x_0)=0\) and \({y_2}'(x_0)=1\). We then have \[y_2(x_0)=C_2x_0=0~~\implies~~C_2=0\text{ or }x_0=0\] Assuming \(C_2\neq0\), this implies the constant \(x_0\) is zero, but let's ignore this for the time being. We also have \[{y_2}'(x_0)=C_2~~\implies~~C_2=1\] and so \(y_2=x\). I'm not quite sure how they're getting \(x-x_0\)...

OpenStudy (idealist10):

So what's the correct answer for this problem?

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

I haven't changed my mind, I'm still not sure how the answer should be \(x-x_0\). Have you tried asking your instructor?

OpenStudy (idealist10):

Yes, but he's not sure either.